OpenStudy (anonymous):

To graph the equation y equals negative one half times x minus 4 using the slope and y-intercept, you would have to do which of the following?

OpenStudy (anonymous):

Graph the y-intercept, 4. From that point, move down 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, -4. From that point, move up 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, -4. From that point, move down 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, 4. From that point, move up 1 and to the right 2 to plot another point. Then draw a line that passes through both points.

OpenStudy (debbieg):

First of all, what's the y-intercept?

OpenStudy (debbieg):

y equals negative one half times x minus 4 Here's the equation: \(\large y=\dfrac{1}{2}x-4\) Where, in that equation, do you find the y-intercept?

OpenStudy (anonymous):

the begining

OpenStudy (debbieg):

I'm not sure what that means. In the equation - what is the y-intercept?

OpenStudy (debbieg):

Your equation is in the form y=mx+b Where in that equation do you find the value of the y-intercept?

OpenStudy (anonymous):

the fraction?

OpenStudy (debbieg):

In a line equation of the form: \(\large y=mx+b\) which piece is the y-intercept? C'mon, I know you've covered this! :)

OpenStudy (anonymous):

i really havent:/

OpenStudy (debbieg):

Look at your equation: \[\large y=\dfrac{1}{2}x-4\] Compare it to general form for the equation of a line: \[\large y=mx+b\]

OpenStudy (debbieg):

Then why do you have these problems? Seriously - I'm not trying to give you a hard time. Help me understand. What class is this? If you are being assigned these problems, then it must be that you are expected to know how to do them.

OpenStudy (anonymous):

its virtual school and i dont unstand any of it. thats why i have to keep taking algebra

OpenStudy (debbieg):

OK, so is this for an algebra class then? Have you covered equations of lines in the class? Or taken the online lesson or whatever? (I don't know much about virtual school, so maybe you can teach me. :)

OpenStudy (anonymous):

its like where they give you a bunch of assignments and you have to complete them all for a credit. the assignments are random but i still have to finish them

OpenStudy (debbieg):

OK, but there has to be a LEARNING component.... some online lesson or lecture on the topic, right?

OpenStudy (anonymous):

i dont have any of that just assignments

OpenStudy (debbieg):

What level of school is this?

OpenStudy (anonymous):

algebra 1

OpenStudy (debbieg):

When you say "just assignments" do you mean just all problems to complete? Isn't there instructional material of some type? I find it hard to believe that you are just handed problems and told to do them without any way to learn the concepts. What am I missing here?

OpenStudy (debbieg):

OK, look. An equation of the form y=mx+b graphs as a line. That means that any point (x,y) that "satisfies" (makes it a true statement) the equation, is a point that will be on the graph.

OpenStudy (debbieg):

b is called the y-intercept because it is the "y coordinate" (do you know what I mean by that?) of the point where the line crosses the y-axis. Every point on the y-axis has x-coordinate of 0, so the equation is satisfied for x=0 if: y=m(0)+b y=b So the point (0,b) is where it crosses the y-axis.

OpenStudy (debbieg):

So what is the y-intercept for your equation here?